eviant:

PLS HELP!!! Let f(x) = -2x - 7 and g(x) = -4x + 6. Find (f*g)(-5)

Shadow:

\[(f \times g)(-5) = (-2(-5) - 7) \times (-4(-5) + 6)\]

Shadow:

From here it is just PEMDAS

Shadow:

@eviant Do you know how to proceed?

eviant:

@Shadow I multiplied the parentheses above and got 78

Shadow:

I got the same.

eviant:

@Shadow that's not an answer choice

eviant:

3 26 –6 –59 ^^^those are the choices

eviant:

@Shadow

Shadow:

Is it f times g or f of g

eviant:

f

Shadow:

Yep, it's not multiplication

Shadow:

\[(f o g)(-5) = (f(g(-5)) = (-2(-4(-5) + 6) - 7)\]

Shadow:

* is a sign sometimes used for multiplication. In these cases, it is best to post a picture if you don't know how to represent the problem.

eviant:

-59?